OpenStudy (anonymous):

find the slope of 2x+y=8

OpenStudy (anonymous):

move 2x to the other side

OpenStudy (anonymous):

y = -2x + 8

OpenStudy (anonymous):

y = mx +b

OpenStudy (anonymous):

m = slope

OpenStudy (anonymous):

m = -2

OpenStudy (anonymous):

slope = -2

OpenStudy (anonymous):

why negative not positive?

jimthompson5910 (jim_thompson5910):

because you are subtracting 2x from both sides (to isolate y)

jimthompson5910 (jim_thompson5910):

2x+y=8 2x+y-2x=8-2x ... Subtract 2x from both sides. y = -2x + 8

jimthompson5910 (jim_thompson5910):

we subtract 2x because on the left side, 2x - 2x turns into 0x or just 0 (and it goes away leaving just y)

OpenStudy (anonymous):

u subtracted the 2x